# Is Causality Necessary for Efficient Portfolios? A Computational Perspective on Predictive Validity and Model Misspecification

### v3

@@ Abstract
A recent line of research has argued that causal factor models are necessary for portfolio optimization, claiming that structurally misspecified models inevitably produce inverted signals and nonviable frontiers. This paper challenges that view. We show, through theoretical analysis, simulation counterexamples, and empirical validation, that predictive models can remain operationally valid even when structurally incorrect. Our contributions are fourfold. First, we distinguish between directional agreement, ranking, and calibration, proving that sign alignment alone does not ensure efficiency when signals are mis-scaled. Second, we establish that structurally misspecified signals can still yield convex and viable efficient frontiers provided they maintain directional alignment with true returns. Third, we derive and empirically confirm a quantitative scaling law that shows how Sharpe ratios contract smoothly with declining alignment, thereby clarifying the role of calibration within the efficient set. Fourth, we validate these results on real financial data, demonstrating that predictive signals, despite structural imperfections, can support coherent frontiers. These findings refine the debate on causality in portfolio modeling. While causal inference remains valuable for interpretability and risk attribution, it is not a prerequisite for optimization efficiency. Ultimately, what matters is the directional fidelity and calibration of predictive signals in relation to their intended use in robust portfolio construction.
@@ 1 Introduction
Recent contributions in empirical finance have argued that portfolio optimization requires causal factor models, contending that structurally misspecified predictors, those that omit relevant variables or misrepresent functional relationships, necessarily generate distorted signals and collapse efficient frontiers (López de Prado and Zoonekynd,, 2024; López de Prado et al.,, 2024, 2025). Within this perspective, causal identifiability is presented as a prerequisite for viability: without it, signal inversion and optimization failure are presumed unavoidable.
@@
This paper challenges the universality of that claim. Through theoretical analysis, controlled simulations, and empirical evidence, we demonstrate that predictive models can remain operationally effective for portfolio construction even when structurally misspecified. In particular, we show that signals derived from such models often preserve directional alignment with the true return vector, which is sufficient to sustain convex and non-degenerate efficient frontiers under classical mean–variance optimization (Almgren and Chriss,, 2005). Inefficiency, in practice, arises more frequently from miscalibration of signal magnitudes rather than from systematic sign inversion.
@@
The contributions of this paper are fourfold. First, we formalize the distinction between directional agreement, ranking, and calibration, proving that directional correctness is necessary but not sufficient for efficiency. Second, we establish geometric results showing that structurally misspecified but directionally aligned signals preserve frontier convexity and optimization validity. Third, we extend the portfolios-from-sorts framework of Almgren and Chriss, (2005) by deriving a scaling law that links Sharpe ratio performance to the cosine alignment between true and surrogate signals. This refinement clarifies that correct ranking ensures membership in the efficient set, while calibration governs quantitative efficiency within it. Fourth, we validate these results using both synthetic data and S&P 500 equity returns, showing that predictive signals lacking causal identifiability can nonetheless support coherent and actionable frontiers.
@@
These findings contribute to the broader debate in finance and econometrics concerning the relative roles of causality and prediction. As emphasized by Shmueli, (2010), explanatory and predictive models serve distinct purposes and should be evaluated against different criteria. Recent critiques have underscored the epistemic limitations of causal reductionism in financial systems: because markets are adaptive and reflexive, unidirectional causal inference may be unstable or conceptually ill-suited (Polakow et al.,, 2024). From this vantage, calibration and directional validity constitute more pragmatic benchmarks for assessing model utility in portfolio design than structural fidelity per se. Related work on model risk and empirical fragility further illustrates the dangers of overstating causal claims in environments characterized by adaptation and self-reference (Harvey and Liu,, 2014; Bailey et al.,, 2017).
@@
We do not dispute the value of causal inference, which remains indispensable for structural attribution, regime monitoring, and stress testing (Wilcox and Gebbie,, 2014; Lo,, 2017). Rather, we caution against overstating its necessity for optimization. What ultimately matters for decision-making is not causal identifiability in the strict sense, but the directional fidelity and calibration of predictive signals relative to their intended application in portfolio construction. In this respect, predictive modeling provides a robust and operationally effective alternative, and often a complement, to causal-first approaches.
@@
The remainder of the paper is structured as follows. Section 2 reviews related work on causal and predictive modeling in finance. Section 3 develops the theoretical framework, establishing results on directionality, ranking, and calibration under misspecification. Section 4 presents empirical validation through simulations and market data. Section 5 examines implications for portfolio theory and modeling practice. Section 6 concludes.
2 Literature Review: Predictive Validity vs. Causal Necessity in Portfolio Modeling
@@
A growing strand of recent literature in asset pricing has emphasized the role of causal factor models as a prerequisite for portfolio efficiency. According to this view, models that omit causal structure, whether due to unobserved confounders, misspecified functional forms, or reliance on purely associational data, are expected to suffer from systematic directional failure, yielding distorted alphas and invalid optimization outcomes (López de Prado et al.,, 2024; López de Prado and Zoonekynd,, 2024; López de Prado et al.,, 2025).
Yet this causal necessity argument rests on assumptions that often prove restrictive in practice. In particular, it underestimates the operational value of predictive modeling, especially in noisy, high-dimensional financial environments where out-of-sample generalization may matter more than structural interpretability. The methodological distinction is well established. Shmueli, (2010) highlight the contrast between explanatory models, which seek to uncover causal relationships, and predictive models, which prioritize forecasting accuracy. Similarly, Breiman, (2001) describes the divide between the “data modeling” culture oriented toward interpretability and the “algorithmic modeling” culture that emphasizes predictive performance.
@@
Finance magnifies this tension. Unlike controlled experimental sciences, financial markets rarely provide stable instruments, suffer from pervasive nonstationarity, and exhibit reflexive dynamics in which causal graphs may be both unstable and unobservable. In such settings, predictive validity frequently dominates causal alignment as the more pragmatic standard of model utility. Moreover, portfolio optimization is notoriously sensitive to estimation error. Seminal work by Best and Grauer, (1991), Chopra and Ziemba, (1993), and Britten-Jones, (1999) demonstrates that even small errors in expected return estimates can induce large swings in portfolio weights. To mitigate this fragility, the literature has developed a series of robustification strategies, including the Black–Litterman model (Black and Litterman,, 1992), constrained optimization approaches (Jagannathan and Ma,, 2003), and shrinkage-based estimators. While Michaud, (1989) introduced a resampling approach where shrinkage was only a byproduct, the more direct lineage of shrinkage estimators originates with the James–Stein estimator (James and Stein,, 1961) and modern covariance shrinkage techniques such as those of Ledoit and Wolf, (2004). Further, DeMiguel et al., (2009) show that simple heuristics, such as equal-weighted portfolios, can outperform optimized allocations when estimation error is severe. Collectively, these results underscore the practical limitations of strictly structural models.
@@
Against this backdrop, predictive-first approaches have gained momentum. Gu et al., (2020) demonstrate that machine learning methods, despite their lack of causal interpretability, deliver superior forecasts and robust portfolio allocations. Similarly, Haugen and Baker, (1996) and Harvey et al., (2016) provide evidence that return-predictive signals can remain operationally effective even without causal grounding. In parallel, researchers have advanced new dependence measures designed to overcome the shortcomings of correlation-based models without appealing to causal structure. For example, Opdyke, (2021) proposes positive definite dependence measures that enable richer associational modeling and improved stress-testing, while Viole and Nawrocki, (2016) develops partial-moment dependence metrics that further broaden the class of non-causal approaches to portfolio design. Together, these contributions demonstrate that non-causal dependence modeling constitutes a viable and expanding alternative to fully specified structural methods.
While causal inference retains value for specific tasks such as risk decomposition or regime identification, the broader evidence indicates that it is not a universal prerequisite for portfolio efficiency. The decisive question is whether a model generates signals that are directionally accurate, properly scaled, and operationally useful for allocation. Under this perspective, predictive validity is often sufficient, and, in many practical settings, more effective, than strict causal specification. The next sections formalize this claim by analyzing the conditions under which predictive signals remain viable despite structural misspecification, and by quantifying how calibration, rather than causal structure, governs realized efficiency.
@@ 3.1.1 Bias from Omitting a Confounder
with $Z$ and $\eta$ independent, so that111(López de Prado et al.,, 2024) incorrectly report $\mathrm{Var}(F_{2})=1+\delta^{2}$, which would be the case only if $\eta\sim\mathcal{N}(0,1)$. With their definition $\eta\sim\mathcal{N}(0,1-\delta^{2})$, the correct result is $\mathrm{Var}(F_{2})=1$. This difference affects the exact expression for the bias in $\hat{\beta}_{n}$: under their assumption $\hat{\beta}_{n}=(\beta_{n}+\gamma_{n}\delta)/(1+\delta^{2})$, whereas under the correct normalization $\hat{\beta}_{n}=\beta_{n}+\gamma_{n}\delta$. In both cases, however, omitting $Z$ induces a $\delta$-dependent bias, so the qualitative conclusion remains unchanged.:
@@
This shows that omitting the confounder $Z$ introduces systematic bias into the estimated factor loading $\hat{\beta}_{n}$, depending on the true confounder exposure $\gamma_{n}$ and its correlation $\delta$ with $F_{2}$ (López de Prado et al.,, 2024).
@@ 3.1.2 Resulting Portfolio Inefficiency
Suppose an investor constructs a minimum-variance portfolio targeting exposure $c^{\top}=[0,1]$ using the misspecified loadings:
$B_{\text{miss}}=\begin{bmatrix}\hat{\beta}_{1}\\ \hat{\beta}_{2}\end{bmatrix}=\begin{bmatrix}\beta_{1}+\gamma_{1}\delta\\ \beta_{2}+\gamma_{2}\delta\end{bmatrix}$
The optimizer solves:
$\omega^{*}_{\text{miss}}=B_{\text{miss}}^{\top}{}^{-1}c^{\top}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\begin{bmatrix}-1\\ 1\end{bmatrix}$
We evaluate the true exposure delivered by this portfolio:
$\omega^{*}_{\text{miss}}{}^{\top}B_{\text{true}}=\left(\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\begin{bmatrix}-1&1\end{bmatrix}\right)\begin{bmatrix}\gamma_{1}&\beta_{1}\\ \gamma_{2}&\beta_{2}\end{bmatrix}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\left(\gamma_{2}-\gamma_{1},\ \beta_{2}-\beta_{1}\right)$
This actual exposure deviates from the intended $[0,1]$ unless $\gamma_{1}=\gamma_{2}=0$, i.e., no confounding. The portfolio is therefore inefficient even though the investor used valid optimization logic on misspecified inputs. In summary, from (López de Prado et al.,, 2024, 2025) even when expected returns and covariances are correctly specified, factor misspecification alone is sufficient to induce portfolio inefficiency. This does not require causal reasoning, it arises purely from statistical misspecification, counter to arguments made in (López de Prado et al.,, 2025).

### v4

@@ Abstract
Portfolio optimization is increasingly argued to require causally identified return predictors to avoid signal inversion and optimization failure. This paper re-examines this claim by studying when predictive signals yield viable efficient frontiers, even under structural misspecification. We show that causal identification is not necessary for portfolio efficiency within static mean–variance and closely related quadratic portfolio optimization frameworks. Instead, efficiency is governed by geometric sufficiency conditions on predictive signals: directional alignment, ranking preservation, and calibration. We formally decompose portfolio efficiency into these three components and show that miscalibration alone attenuates Sharpe ratios even when alignment and ranking are preserved. Robustness is characterized as smooth degradation rather than collapse, with explicit attenuation behavior and continuity of performance under increasing misspecification. The theoretical results are supported by simulations and empirical analysis. Empirical validation combines equity-based illustrations with a large global bond universe spanning multiple currencies, countries, sectors, maturities along the term structure, seniority classes, and credit ratings, together with high-dimensional stress tests, nonlinear data-generating processes, rolling-window analyses, covariance regularization, realistic portfolio constraints, and bootstrap-based statistical validation. Across these settings, optimization geometry remains well behaved whenever directional alignment is preserved. The results clarify the boundary between causality and portfolio optimization: causality may inform signal representation, but portfolio efficiency at the optimization stage is a geometric property conditional on a given representation.
@@ 1 Introduction
Recent contributions in empirical finance have argued that portfolio optimization requires causal factor models, contending that structurally misspecified predictors (those that omit relevant variables or misrepresent functional relationships) necessarily generate distorted signals and collapse efficient frontiers. Within this perspective, causal identifiability is presented as a prerequisite for viability: without it, signal inversion and optimization failure are presumed unavoidable (López de Prado et al., 2025a).
A recurring claim in the recent causal factor–investing literature is that improvements in predictive performance achieved through causal identification are necessary, or at least taken to be strongly diagnostic, for portfolio efficiency. This claim conflates two distinct questions. From the perspective of modern portfolio theory, efficiency is not a property of individual predictors, but of the joint geometry induced by the return signal and the risk model. Conditional on a given admissible risk–return geometry, efficiency is governed by how predictive signals align, rank, and scale within that space, rather than by whether those signals admit a causal interpretation. Causal structure may improve signal construction, but it is not intrinsic to the definition or feasibility of efficient portfolios within this framework. Throughout the paper, by quadratic portfolio optimization frameworks, we refer to static mean–variance optimization and equivalent formulations that admit a quadratic risk objective.
@@
This paper challenges the universality of the causal-necessity claim. Through theoretical analysis, controlled simulations, and empirical evidence, we demonstrate that predictive models can remain operationally effective for portfolio construction even when structurally misspecified. In particular, we show that signals derived from such models often preserve directional alignment with the true return vector, which is sufficient to sustain convex and non-degenerate efficient frontiers under classical mean–variance optimization. Inefficiency, in practice, arises more frequently from miscalibration of signal magnitudes than from systematic sign inversion.
More generally, the results of this paper establish that causal identifiability is not a necessary condition for portfolio efficiency within the class of quadratic portfolio optimization problems studied here. Portfolio diversification is inherently a geometric problem: efficiency is determined by the relative orientation, ordering, and scaling of predictive signals within a given risk–return geometry. As shown in this work, three geometric properties—directional alignment, ranking preservation, and calibration—jointly characterize investment efficiency independently of the mechanism generating the signals.
Two clarifications are therefore essential to properly situate this contribution. First, the claim advanced here is not that causality is irrelevant, but that it is not intrinsic to the objective of static portfolio optimization. In the classical formulation of modern portfolio theory (Markowitz, 1952), efficiency is defined geometrically through the joint distribution of returns and their covariance structure, rather than through the causal mechanisms generating those returns. Treating causal identifiability as a prerequisite for efficiency therefore conflates explanatory objectives with optimization geometry.
Second, when causality is incorporated into portfolio construction, it enters through the specification of the diversification geometry itself rather than as a requirement imposed on return prediction. Related work by Rodriguez Dominguez (2023, 2025) addresses a complementary setting in which diversification is modeled through time-varying sensitivities to common causal drivers. In that framework, alignment, ranking, and calibration are evaluated relative to the induced local geometry, illustrating that even causal formulations ultimately operate through geometric conditions at the optimization stage.
@@
The contributions of this paper are fourfold. First, we introduce and formalize a three-way distinction between directional alignment, ranking preservation, and calibration in portfolio optimization, showing that directional correctness is necessary but not sufficient for portfolio efficiency. Second, we establish geometric results demonstrating that structurally misspecified yet directionally aligned signals preserve frontier convexity and optimization viability. Third, we extend the portfolios-from-sorts framework of Almgren and Chriss (2005) by explicitly separating its ranking-based insight from other dimensions of predictive validity and by deriving a scaling law that links Sharpe ratio performance to the cosine alignment between true and surrogate signals. While Almgren and Chriss (2005) shows that correct ranking ensures membership in the efficient set, we further establish that calibration governs quantitative efficiency within that set. Fourth, we validate these results using both synthetic data and S&P 500 equity returns, showing that predictive signals lacking causal identifiability can nonetheless support coherent and actionable efficient frontiers.
@@
These findings contribute to the broader debate in finance and econometrics concerning the relative roles of causality and prediction. As emphasized by Shmueli (2010), explanatory and predictive models serve distinct purposes and should be evaluated against different criteria. Recent critiques have underscored the epistemic limitations of causal reductionism in financial systems: because markets are adaptive and reflexive, unidirectional causal inference may be unstable or conceptually ill-suited (Polakow et al., 2024). From this vantage, calibration and directional validity constitute more pragmatic benchmarks for assessing model utility in portfolio design than structural fidelity per se.
@@
We do not dispute the value of causal inference, which remains indispensable for structural attribution, regime monitoring, and stress testing (Wilcox and Gebbie, 2014; Lo, 2017). Rather, we caution against overstating its necessity for optimization. What ultimately matters for portfolio construction is not causal identifiability per se, but the directional fidelity, ranking stability, and calibration of predictive signals relative to the chosen diversification geometry.
To clarify the logic of the argument and the scope of the results, Figure 1 provides a structured overview of how predictive models, geometric signal properties, and optimization interact in this paper.
Recent arguments for the necessity of causal factor modeling rely on a small number of distinct claims regarding misspecification, signal inversion, optimization failure, and fragility of efficiency under noise. For clarity, these claims and their logical refutations are summarized in Table 3 in Appendix B. The main text addresses them sequentially: Section 3.1 examines bias under omitted variables, Section 3.2 refutes the generic inversion claim, Sections 3.3 and 3.4 characterize the conditions under which optimization geometry remains viable, and Section 3.5 establishes continuity and robustness under approximate cancellation.
@@
The remainder of the paper is structured as follows. Section 2 reviews related work on causal and predictive modeling in finance. Section 3 develops the theoretical framework. Section 4 presents empirical validation through simulations and market data. Section 5 examines implications for portfolio theory and modeling practice. Section 6 concludes.
Predictive Modellinear / nonlinearcausal or non-causalStructural Misspecificationomitted variablesnonlinear confoundingSignal Geometrydirectional alignmentranking preservationcalibrationFailure Modesattenuation (miscalibration)inversion (special case)Geometric Robustnessconvexity preservedsmooth degradationPortfolio Optimizationmean–variance geometryquadratic riskEfficient Frontierfeasible / convexSharpe scales with alignmentEmpirical Validationsimulations + S&P 500high-dimensional stress tests
Figure 1: Logical structure of the framework and the paper. Predictive models—whether causal or non-causal—produce surrogate signals whose geometric properties determine optimization behavior. Structural misspecification primarily induces attenuation through miscalibration rather than generic inversion. Portfolio efficiency is governed by alignment, ranking, and calibration within a given risk–return geometry, with robustness characterized by smooth degradation rather than frontier collapse.
2 Literature Review
2.1 Causal Identification and Causal Representations in Finance
@@
A growing strand of recent literature in asset pricing has emphasized the role of causal factor models in portfolio construction. According to this view, models that omit causal structure (whether due to unobserved confounders, misspecified functional forms, or reliance on purely associational data) are argued to risk distorted inference and unstable optimization outcomes (López de Prado et al., 2025a). This perspective is further developed in López de Prado (2023); López de Prado and Zoonekynd (2026); López de Prado et al. (2025b), where causal identification is presented as a safeguard against factor mirages and misleading allocation decisions.
More recently, this line of work has been extended through causal network and causal discovery approaches applied to asset pricing and factor investing. Examples include causal spillover networks across asset and factor systems (Konstantinov and Fabozzi, 2026), causal network representations for factor investing (Howard et al., 2025), and automated causal discovery in financial markets under nonstationarity (Sokolov et al., 2024; Sadeghi et al., 2025; Sadeghi and Simonian, 2025). Related contributions formalize causal inference frameworks for asset pricing (Haddad et al., 2025) and survey the rapidly expanding use of causal methods across finance and insurance (Kumar et al., 2025). Several of these works implicitly or explicitly suggest that causal identification is required to avoid systematic portfolio inefficiency, particularly in factor-based investment settings.
2.2 Prediction, Estimation Error, and Portfolio Optimization
Yet causal-necessity arguments rest on assumptions that are often restrictive in practice. In particular, they conflate explanatory validity with the requirements of portfolio efficiency. From the standpoint of modern portfolio theory, efficiency is fundamentally a diversification concept, defined relative to the geometry induced by the asset universe and the admissible risk model, rather than by the causal provenance of return predictors.
The methodological distinction between explanation and prediction is well established. Shmueli (2010) emphasize the contrast between explanatory models, which seek causal interpretation, and predictive models, which prioritize forecasting accuracy. Similarly, Breiman (2001) describes the divide between the ”data modeling” culture oriented toward interpretability and the ”algorithmic modeling” culture emphasizing predictive performance. Finance magnifies this tension. Financial markets rarely provide stable instruments, exhibit pervasive nonstationarity, and display reflexive dynamics in which causal graphs may be unstable or unobservable. In such environments, predictive validity frequently dominates causal alignment as the pragmatic standard of model utility.
@@ 2.2 Prediction, Estimation Error, and Portfolio Optimization
Moreover, portfolio optimization is notoriously sensitive to estimation error. Seminal work by Best and Grauer (1991), Chopra and Ziemba (1993), and Britten-Jones (1999) shows that even small errors in expected return estimates can induce large swings in portfolio weights. This fragility has motivated a wide range of robustification strategies, including the Black–Litterman model (Black and Litterman, 1992), constrained optimization (Jagannathan and Ma, 2003), and shrinkage-based estimators (Ledoit and Wolf, 2004). Further, DeMiguel et al. (2009) demonstrate that simple heuristics, such as equal-weighted portfolios, can outperform optimized allocations when estimation error is severe.
@@
Against this backdrop, predictive-first methodologies have gained momentum. Gu et al. (2020) show that machine learning models, despite limited causal interpretability, deliver strong predictive performance and competitive portfolio allocations. Related evidence suggests that return-predictive signals can remain operationally effective even without explicit causal grounding (Haugen and Baker, 1996; Harvey et al., 2016). In parallel, a growing literature has explored non-causal dependence measures that extend beyond correlation without invoking causal structure (Opdyke, 2021; Viole and Nawrocki, 2016). Together, these contributions demonstrate that associational and predictive modeling constitutes a viable alternative to fully specified structural approaches.
2.3 Scope of the Critique and the Necessity Question
Recent causal discovery and causal network studies primarily aim to improve interpretability, stress testing, or predictive stability rather than to analyze portfolio efficiency directly. While such methods may inform signal construction or risk attribution, they do not characterize the conditions under which efficient frontiers remain viable within a given optimization framework. In that setting, portfolio efficiency depends on how signals interact within a diversification geometry, rather than on the correctness or completeness of an underlying causal graph.
To clarify how the present contribution relates to prior work, it is useful to distinguish three broad strands in the literature. A first strand establishes that directional or ranking information may suffice for portfolio efficiency under stylized or correctly specified settings (e.g., Almgren and Chriss (2005)). A second strand documents empirically that non-causal or purely predictive signals, including machine-learning-based forecasts, can yield competitive portfolios in practice (e.g., DeMiguel et al. (2009); Gu et al. (2020)). A third strand emphasizes the fragility of portfolio optimization under estimation error or misspecification, often motivating structurally interpretable or causal modeling approaches (e.g., Best and Grauer (1991); Chopra and Ziemba (1993)).
None of these strands, however, directly addresses the necessity question that motivates the present study: whether causal identification is required for portfolio efficiency within the static mean–variance and related optimization frameworks considered in this paper, and if not, what minimal conditions replace it. The contribution of this paper is to resolve this question by formalizing geometric sufficiency conditions based jointly on directional alignment, ranking preservation, and calibration. We show that these conditions govern portfolio efficiency independently of causal identification within this framework, persist under nonlinear data-generating processes, and degrade continuously rather than collapsing under misspecification. Accordingly, the analysis that follows is not concerned with proposing new predictors or causal discovery methods, but with clarifying the boundary between causal relevance and predictive sufficiency in portfolio optimization.
@@ 3 Methods
This section develops the theoretical analysis around three claims that underlie recent arguments for the necessity of causal modeling in portfolio optimization: (i) that structural misspecification generically induces signal inversion; (ii) that portfolio inefficiency arises only through such inversion; and (iii) that efficient frontiers collapse in the absence of causal identification. All results in this section are derived within static mean–variance and related quadratic portfolio optimization frameworks. Sections 3.13.2 analyze the behavior of predictive signals under omitted variables and confounding, while Sections 3.33.5 characterize the conditions under which portfolio optimization remains well posed within these settings.
While individual elements such as directional information, ranking, and estimation error sensitivity have appeared separately in prior portfolio and asset pricing literature, the contribution of this paper is to resolve an explicit necessity question that has remained implicit: whether causal identification is required for portfolio efficiency within the class of optimization frameworks considered here. We provide new results that jointly characterize efficient frontier viability under misspecification through geometric sufficiency conditions, identify calibration as the quantitative driver of Sharpe attenuation even under correct alignment and ranking, and show that inefficiency degrades continuously—rather than collapsing—under nonlinear and high-dimensional misspecification. These results are not implied by existing work that considers alignment, ranking, or estimation error in isolation.
@@ 3.1 Preliminary Work: Analyzing Misspecification in Linear Factor Models
Notation. Throughout Section 3, $\mu$ denotes the true expected return vector, $\tilde{\mu}$ a surrogate or misspecified signal, $\Sigma$ the return covariance matrix, and $\omega$ portfolio weights. All inner products and norms are defined with respect to the geometry induced by $\Sigma^{-1}$ unless stated otherwise.
@@ 3.1.1 Bias from Omitting a Confounder
with $Z$ and $\eta$ independent, so that111(López de Prado et al., 2024) compute this variance under an alternative normalization, reporting $\mathrm{Var}(F_{2})=1+\delta^{2}$, which would correspond to the specification $\eta\sim\mathcal{N}(0,1)$. Under the definition stated in their model, $\eta\sim\mathcal{N}(0,1-\delta^{2})$, the correct result is $\mathrm{Var}(F_{2})=1$. This difference affects the scaling of the omitted-variable bias in $\hat{\beta}_{n}$: under their normalization $\hat{\beta}_{n}=(\beta_{n}+\gamma_{n}\delta)/(1+\delta^{2})$, whereas under the stated normalization $\hat{\beta}_{n}=\beta_{n}+\gamma_{n}\delta$. In both cases, however, omitting $Z$ induces a $\delta$-dependent bias, so the qualitative conclusion regarding misspecification remains unchanged.:
@@
This shows that omitting the confounder $Z$ introduces systematic bias into the estimated factor loading $\hat{\beta}_{n}$, depending on the true confounder exposure $\gamma_{n}$ and its correlation $\delta$ with $F_{2}$ (López de Prado et al., 2024).
@@ 3.1.2 Resulting Portfolio Inefficiency
Suppose an investor constructs a minimum-variance portfolio targeting exposure $c^{\top}=[0,\,1]$ using the misspecified loadings
$B_{\text{miss}}=\begin{pmatrix}\hat{\beta}_{1}\\ \hat{\beta}_{2}\end{pmatrix}=\begin{pmatrix}\beta_{1}+\gamma_{1}\delta\\ \beta_{2}+\gamma_{2}\delta\end{pmatrix}$,
as in López de Prado et al. (2024, 2025a). The optimizer solves [^2] This expression implicitly assumes a two-asset setting in which the exposure constraint admits a one-dimensional null space. While not stated explicitly in (López de Prado et al., 2024, 2025a), this restriction is required for the above expression to be well defined.
$\omega^{\ast}_{\text{miss}}=B_{\text{miss}}^{\top-1}c^{\top}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\begin{pmatrix}-1\\ \;\,1\end{pmatrix}$.
We evaluate the true exposure delivered by this portfolio under the correctly specified loading matrix
$B_{\text{true}}=\begin{pmatrix}\gamma_{1}&\beta_{1}\\ \gamma_{2}&\beta_{2}\end{pmatrix}$.
A direct calculation gives
$\omega^{\ast\top}_{\text{miss}}B_{\text{true}}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\left(\gamma_{2}-\gamma_{1},\;\beta_{2}-\beta_{1}\right).$
Unless $\gamma_{1}=\gamma_{2}=0$, that is, unless the omitted confounder has identical exposure across assets, the realized exposure deviates from the intended target [0, 1][0,\,1]. Thus, even within the work of López de Prado et al. (2024, 2025a), factor misspecification alone suffices to induce portfolio inefficiency, despite the use of a formally valid optimization procedure. This result follows purely from statistical misspecification of the factor loadings and does not rely on causal interpretation. In particular, inefficiency arises without requiring signal inversion, a distinction that will be central in the subsequent analysis.